OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5).

OpenStudy (anonymous):

same thing. 52

OpenStudy (anonymous):

im guessing thats an x^2?

OpenStudy (anonymous):

Yes it is.

OpenStudy (lalaly):

\[2(5)^2+2 = 50+2=52\]

OpenStudy (anonymous):

yup, just put 5 wherever you see x in the equation and solve it. \[f(x) = 2x^2+2\]\[f(5)=2*(5)^2+2\]\[f(5)=52\]